A solid sphere of radius R is placed at rest on a rough horizontal surface. The sphere is given a horizontal impulse at height h above its lowest point. Consequently, it starts rolling without slipping. The value of h is

Grade:11

2 Answers

Vikas TU
7 years ago
Moment zero at point of contct.
Hence, F*h = I*a/R
Fh = (2mR^2/5 + mR^2)a/R
h = (7mR/5)a meter
where a is the accleration and m is the mass of the sphere.
ruhi
5 years ago
according to the question the sphere was is in rest therefore the net torque=0
I×h - F×R =0
MR^2 ×h =(2/5 )MR^2 ×R
h/R= 2/5
this is the required ratio...
here R= radius of the sphere
I= moment of inertia applied to the sphere
F= moment of inertia about diameter